OpenStudy (anonymous):

help please anyone good in calculus and please explain step by step it would be really great.

OpenStudy (anonymous):

A consulting firm conducts training sessions for employees of various companies. The charge to a company sending employees to a session is $50 per employee, less $0.50 for each employee in excess of 10. That is, for example, if twelve employees are sent, the charge per employee would be $49.00 and the total prorated charge to the company would be 12(49.00)=$558.00. The consulting firm further has a fixed total charge for groups of x or more where x is the number that maximizes the prorated group charge. What should x be, and what is the maximum total group charge to the company?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (dumbcow):

first write equation for charge in terms of x: \[C = x[50 - 0.5(x-10)] = x(55 - 0.5x)\] \[C = 55x - 0.5x^{2}\] we want to maximize charge so set derivative equal to zero \[\frac{dC}{dx} = 55 - x = 0\] \[x = 55\] plug it into C equation to get max charge